1. If 1.3 L of 3.0M potassium hydroxide completely neutralizes 0.7 L of carbonic acid




(H2CO3), what is the concentration of the carbonic acid? 2KOH+H,CO, K,CO,+2H₂O




2. How many moles of HCl are present in 0.07 L of 6 M HCI? 3. I add 7 mL of 2 M NaOH to 50 mL of water. What is the new concentration?




Rates of Reactions-Le Chatlier's Principle




2KOH+H₂CO,→K.CO,+2H2O overall energy of reaction (AH)=-683.84 kJ/mol 1. Describe 2 ways you could increase the amount of products of this reaction at




equilibrium. 2. Describe 2 ways you could increase the amount of reactants of the reaction at




equilibrium.

1
Expert's answer
2022-06-05T21:09:04-0400

1.

n (KOH)= CM * V = 1,3 * 3 = 3,9 mol


2KOH + H2CO3 = K2CO3 + 2H2O


2mol KOH — 1mol H2CO3

3,9mol KOH — X


X = 3,9 * 1 / 2 = 1,95 mol H2CO3


CM(H2CO3) = n / V = 1,95 / 0,7 = 2,786 mol/l

Answer: 2,786 mol/l H2CO3


2.

n(HCl)= CM * V = 6 * 0,07 = 0,42 mol

Answer: 0,42 mol HCl


3.

CM (old) = 2 mol/l

V (old) = 7 ml = 0,007 l

V (new) = 57 ml = 0,057 l


n(NaOH)= CM * V = 2 * 0,007 = 0,014 mol


CM(new) = n / V(new) = 0,014 / 0,057 = =0,2456 mol/l

Answer: 0,2456 M NaOH


2KOH + H2CO3 = K2CO3 + 2H2O + 683,84 kJ

1.

Increasing the reaction products is to shift the equilibrium of the chemical reaction to the right. For this:

a) Since the reaction is exothermic, it must be cooled. 

b) water or K2CO3 should be removed from the reaction. 

c) KOH or H2CO3 should be added.


2.

increasing the amount of reactants shifts the equilibrium of this chemical reaction to the left. For this:

a) Since the reaction is exothermic, it must be heating. 

b) K2CO3 or H2O should be added.